# An operator learning perspective on parameter-to-observable maps

### v1

@@ C.3. Proofs for full-field learning of factorized linear functionals
Let $\nu^{\prime}$ satisfy Assumption (A-I). Let $q$ be a linear functional and $L$ be a linear operator such that $L=\sum_{j=1}^{\infty}l_{j}\varphi_{j}\otimes\varphi_{j}$ for eigenbasis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\Sigma^{\prime}=\operatorname{Cov}(\nu^{\prime})$ and eigenvalue sequence $\{l_{j}\}_{j\in\mathbb{N}}\subset\mathbb{R}$. Write $\{\sigma_{j}^{\prime}\}_{j\in\mathbb{N}}$ for the eigenvalues of Σsuperscriptnormal-$\Sigma^{\prime}$. If $q\circ L$ is continuous, then
@@
Lemma C.16 will be used in the proofs of following two theorems. The arguments rely on [16, Theorem 3.9, pp. 18–19].
@@
The supremum is finite because its argument is summable due to $q^{\dagger}\in\mathcal{H}^{r}$. The asserted result follows from [16, Theorem 3.9, pp. 18–19] by using the result for the posterior mean, choosing $\lvert\vartheta_{j}^{\prime}\rvert^{2}=j^{-2(\alpha^{\prime}+r)}$ (i.e, replacing $\alpha^{\prime}$ with $\alpha^{\prime}+r$), and choosing $\delta$ to be a sufficiently small constant. ∎
The theorem for power law QoI decay has a proof similar to the previous one. See 4.8
@@
Replacing $\alpha^{\prime}$ with $\alpha^{\prime}+r+1/2$ in [16, Thm. 3.9, pp. 18–19] completes the proof. ∎
@@ C.4. Proof of sample complexity comparison
if $2s-2r-1-2\beta\leq 0$. This proves the claim. For the end-to-end bound, choose the maximal $s=\beta+r+1/2$ in Theorem 4.7. With optimal $p=s+1/2$, the assumption $p>1/2$ gives $s>0$ so that $\beta+r+1/2>0$ as hypothesized. This condition also satisfies the continuity requirement (A-V) by a similar calculation. The condition $\alpha+s>1$ from Assumption 4.4 is the same as $\alpha+\beta+r>1/2$. Finally, to satisfy the condition $\min(\alpha,\alpha+r+1/2)+\beta>0$ from Theorem 4.8, it suffices for $\alpha+\beta>0$ because the other case has $\alpha+r+1/2+\beta>\max(\alpha,1)>0$. With a common set of hypotheses identified, the convergence rates may now simply be read off from Theorem 4.7 and Theorem 4.8 after plugging in $\alpha^{\prime}=\alpha$ and $s=\beta+r+1/2$. The fact that these bounds hold simultaneously on an event with the asserted probability follows by a union bound and enlarging constants. ∎
@@ C.5. Technical lemmas
We conclude Appendix C with several supporting lemmas. The following two technical results emphasize the non-asymptotic nature of analogous asymptotic bounds on parametrized series from [29, Lemmas 8.1–8.2, pp. 2653–2655]. The first result is useful for controlling the bias error term arising in Appendix C.2.2.
@@
The asserted result may be extracted from the proof of [29, Lemma 8.1]. ∎
@@
Since $j>J_{\mu}\geq\mu^{-1/u}$, it holds that $1\leq 1+\mu^{-1}j^{-u}\leq 2$. Hence, the right hand side of the preceding display is bounded above and below by a constant times $\sum_{j>J_{\mu}}j^{-u}$. By comparison to an integral as in the proof of Lemma C.18, we obtain
@@
$E\coloneqq\mathopen{}\mathclose{{}\left\{\big{\lVert}\mathcal{C}_{\mu}^{-1/2}(\widehat{\mathcal{C}}-\mathcal{C})\mathcal{C}_{\mu}^{-1/2}\big{\rVert}_{\mathcal{L}(H)}\leq 1/2}\right\}$ (107)
satisfies $\operatorname{\mathbb{P}}(E)\geq 1-e^{-cN}$.
@@
By [26, Lemma 5, p. 13], there exists $c_{0}\in(0,1)$ such that if $\operatorname{tr}(\mathcal{C}_{\mu}^{-1}\mathcal{C})\leq c_{0}N$ and the pushforward $(\Lambda^{1/2})_{\sharp}\nu$ is strongly-subgaussian, then $\operatorname{\mathbb{P}}(E)\geq 1-e^{-c_{0}N}$. An $H$-valued random variable $Z$ (equivalently, its law) is strongly-subgaussian if
@@
for some absolute constant $c^{\prime}>0$ [74, p. 28] and any $\lambda\in\mathbb{R}$. We used independence of the $\{z_{j}\}_{j\in\mathbb{N}}$ to obtain the second equality in the preceding display and subgaussianity to obtain the first inequality. Again by [74, p. 28], $m\langle h,\Sigma h\rangle^{1/2}$ is equivalent to the subgaussian norm of $\langle u,h\rangle$. Thus, $\nu$ is strongly-subgaussian as required.
@@
Following [11, Section 5.2, p. 350] and [20, Section 6.2, p. 26], we recall the following identity for regularized inverses of linear operators. This result is especially useful in conjunction with the cyclic property of the trace for deriving trace bounds.

### v2

@@ C.3. Proofs for full-field learning of factorized linear functionals
Let $\nu^{\prime}$ satisfy Assumption (A-I). Let $q$ be a linear functional and $L$ be a linear operator such that $L=\sum_{j=1}^{\infty}l_{j}\varphi_{j}\otimes\varphi_{j}$ for eigenbasis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ of $\Sigma^{\prime}=\operatorname{Cov}(\nu^{\prime})$ and eigenvalue sequence $\{l_{j}\}_{j\in\mathbb{N}}\subset\mathbb{R}$. Write $\{\sigma_{j}^{\prime}\}_{j\in\mathbb{N}}$ for the eigenvalues of $\Sigma^{\prime}$. If $q\circ L$ is continuous, then
@@
Lemma C.16 will be used in the proofs of following two theorems. The arguments rely on [17, Theorem 3.9, pp. 18–19].
@@
The supremum is finite because its argument is summable due to $q^{\dagger}\in\mathcal{H}^{r}$. The asserted result follows from [17, Theorem 3.9, pp. 18–19] by using the result for the posterior mean, choosing $\lvert\vartheta_{j}^{\prime}\rvert^{2}=j^{-2(\alpha^{\prime}+r)}$ (i.e, replacing $\alpha^{\prime}$ with $\alpha^{\prime}+r$), and choosing $\delta$ to be a sufficiently small constant. ∎
The theorem for power law QoI decay has a proof similar to the previous one. See 4.9
@@
Replacing $\alpha^{\prime}$ with $\alpha^{\prime}+r+1/2$ in [17, Theorem 3.9, pp. 18–19] completes the proof. ∎
@@ C.4. Proof of sample complexity comparison
if $2s-2r-1-2\beta\leq 0$. This proves the claim. For the end-to-end bound, choose the maximal $s=\beta+r+1/2$ in Theorem 4.8. With optimal $p=s+1/2$, the assumption $p>1/2$ gives $s>0$ so that $\beta+r+1/2>0$ as hypothesized. This condition also satisfies the continuity requirement (A-V) by a similar calculation. The condition $\alpha+s>1$ from Assumption 4.5 is the same as $\alpha+\beta+r>1/2$. Finally, since $\nu=\nu^{\prime}$ implies $\alpha=\alpha^{\prime}$, to satisfy the condition $\min(\alpha,\alpha+r+1/2)+\beta>0$ from Theorem 4.9, it suffices for $\alpha+\beta>0$ because the other case has $\alpha+r+1/2+\beta>\max(\alpha,1)>0$. With a common set of hypotheses identified, the convergence rates may now simply be read off from Theorem 4.8 and Theorem 4.9 after plugging in $\alpha^{\prime}=\alpha$ and $s=\beta+r+1/2$. The fact that these bounds hold simultaneously on an event with the asserted probability follows by a union bound and enlarging constants. ∎
@@ C.5. Technical lemmas
We conclude Appendix C with several supporting lemmas. The following two technical results emphasize the non-asymptotic nature of analogous asymptotic bounds on parametrized series from [34, Lemmas 8.1–8.2, pp. 2653–2655]. The first result is useful for controlling the bias error term arising in Appendix C.2.2.
@@
The asserted result may be extracted from the proof of [34, Lemma 8.1]. ∎
@@
Since $j>J_{\mu}\geq\mu^{-1/u}$, it holds that $1\leq 1+\mu^{-1}j^{-u}\leq 2$. Hence, the right-hand side of the preceding display is bounded above and below by a constant times $\sum_{j>J_{\mu}}j^{-u}$. By comparison to an integral as in the proof of Lemma C.18, we obtain
@@
$\mathsf{E}\coloneqq\mathopen{}\mathclose{{}\left\{\big{\lVert}\mathcal{C}_{\mu}^{-1/2}(\widehat{\mathcal{C}}-\mathcal{C})\mathcal{C}_{\mu}^{-1/2}\big{\rVert}_{\mathcal{L}(H)}\leq 1/2}\right\}$ (112)
satisfies $\operatorname{\mathbb{P}}(\mathsf{E})\geq 1-e^{-cN}$.
@@
By [30, Lemma 5, p. 13], there exists $c_{0}\in(0,1)$ such that if $\operatorname{tr}(\mathcal{C}_{\mu}^{-1}\mathcal{C})\leq c_{0}N$ and the pushforward $(\Lambda^{1/2})_{\sharp}\nu$ is strongly-subgaussian, then $\operatorname{\mathbb{P}}(\mathsf{E})\geq 1-e^{-c_{0}N}$. An $H$-valued random variable $Z$ (equivalently, its law) is strongly-subgaussian if
@@
for some absolute constant $c^{\prime}>0$ [86, p. 28] and any $\lambda\in\mathbb{R}$. We used independence of the $\{z_{j}\}_{j\in\mathbb{N}}$ to obtain the second equality in the preceding display and subgaussianity to obtain the first inequality. Again by [86, p. 28], $m\langle h,\Sigma h\rangle^{1/2}$ is equivalent to the subgaussian norm of $\langle u,h\rangle$. Thus, $\nu$ is strongly-subgaussian as required.
@@
As discussed in Remark 4.4, we are able to weaken the strongly-subgaussian requirement on the input training distribution $\nu$ (11) by only requiring its KL expansion coefficients (22) to be pairwise uncorrelated instead of statistically independent. However, the following result shows that this improvement to Assumption 4.3 leads to a strictly worse failure probability for the good event $\mathsf{E}$ (112). Indeed, Lemma C.20 gives $\operatorname{\mathbb{P}}(\mathsf{E}^{\textsf{c}})\leq e^{-cN}$, while the next lemma yields $\operatorname{\mathbb{P}}(\mathsf{E}^{\textsf{c}})\leq 2e^{-cN^{r}}$ with $r=(\alpha+p-1)/(\alpha+p)<1$ being strictly smaller than one. The proof of this result closely mimics the argument in the proof of Lemma C.9 and may be found in [62, Lemma D.26, p. 264–265]
Lemma C.21 (good set without independent KL coefficients).
Let $\mu=\gamma^{2}/N$ and $\mathcal{C}$, $\mathcal{C}_{\mu}$, and $\widehat{\mathcal{C}}$ be as in (75) and (77). Let Assumption 4.2 hold. Suppose that the hypotheses of Assumption 4.3 hold, but instead of the requirement that the normalized KL expansion coefficients $\{z_{j}\}_{j\in\mathbb{N}}$ (22) of $u\sim\nu$ are independent, assume now that the $\{z_{j}\}_{j\in\mathbb{N}}$ are only pairwise uncorrelated. Then there exist constants $c\in(0,1)$ and $N_{0}\geq 1$ (depending only on $\nu$, $\Lambda$, and $\gamma^{2}$) such that if $N\geq N_{0}$, then the event
$\mathsf{E}\coloneqq\mathopen{}\mathclose{{}\left\{\big{\lVert}\mathcal{C}_{\mu}^{-1/2}(\widehat{\mathcal{C}}-\mathcal{C})\mathcal{C}_{\mu}^{-1/2}\big{\rVert}_{\mathcal{L}(H)}\leq 1/2}\right\}$ (114)
satisfies $\operatorname{\mathbb{P}}(\mathsf{E})\geq 1-2\exp(-cN^{\frac{\alpha+p-1}{\alpha+p}})$, where $\alpha$ and $p$ are as in (A2) and (A3).
Following [12, Section 5.2, p. 350] and [22, Section 6.2, p. 26], we recall the following identity for regularized inverses of linear operators. This result is especially useful in conjunction with the cyclic property of the trace for deriving trace bounds.
@@
Acknowledgments
The first author is supported by the high-performance computing platform of Peking University. The second author acknowledges support from the National Science Foundation Graduate Research Fellowship Program under award number DGE-1745301 and from the Amazon/Caltech AI4Science Fellowship, and partial support from the Air Force Office of Scientific Research under MURI award number FA9550-20-1-0358 (Machine Learning and Physics-Based Modeling and Simulation). The third author is supported by the Department of Energy Computational Science Graduate Fellowship under award number DE-SC00211. The second and third authors are also grateful for partial support from the Department of Defense Vannevar Bush Faculty Fellowship held by Andrew M. Stuart under Office of Naval Research award number N00014-22-1-2790. The computations presented in this paper were partially conducted on the Resnick High Performance Computing Center, a facility supported by the Resnick Sustainability Institute at the California Institute of Technology. The authors thank Kaushik Bhattacharya for useful discussions about learning functionals, Andrew Stuart for helpful remarks about the universal approximation theory, and Zachary Morrow for providing the code for the advection–diffusion equation solver.